OpenStudy (anonymous):

Jordan is twice the age of Luke who is 5 years older than Erin. If the total of their ages is 59 how old is Erin?

OpenStudy (anonymous):

defining variables Let "j" be the age of jordan let "l" be the age of luke let "e" be the age of erin jordan is twice the age of luke meaning j= 2l Luke is 5 years older than Erin meaning l=e+5 we can substitute to get j=2(e+5) the total of their ages is 59 meaning l+j+e = 59 substitute for l and j and then solve for e( age of erin)